OpenStudy (anonymous):

f the sequence is geometric, give its common ratio. 800, 400, 200, 100 ... Answer 2 ½ -2 not geometric

OpenStudy (anonymous):

@Shadowys

OpenStudy (anonymous):

@onmng314

OpenStudy (anonymous):

It 2

OpenStudy (anonymous):

1/2 Can you think of why?

OpenStudy (anonymous):

nope? @Shadowys yours

OpenStudy (anonymous):

It decreases by a factor of 2...so it's 1/2... It's the other way around.

OpenStudy (anonymous):

800(1/2)=400 400(1/2)=200 200(1/2=100 etc

OpenStudy (anonymous):

I got it now guys.